A plastic ball is dropped from height of 1m and rebounds several times from floor.if 1.3s elapses from the moment it is dropped to the second impact with the floor, what is the coefficient of restitution.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
For restitution coeff.,
U = root(2gh) => root(20) = > 4.47 m/s
V = root(2gh) + g*1.3 = > 4.47 + 13 => 17.47
 
e = V/U => 17.47/4.47 = > 3.90